Based on a sample of 450 AUS students, we found the average score is 165, with a standard deviation of 6.
a) At least how many percent of students scored between 150 to 180.
b) Assuming the histogram of GPA has a right skewed distribution, calculate the range of GPA that will cover at least 71% of the students?
c) Assuming the histogram of GPA follows the normal distribution, approximately how many students score between 147 to 159?
1
Expert's answer
2020-12-21T17:23:49-0500

Let "X=" the score of the student: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0,1)"

Given "\mu=165, \sigma=6"

a)


"P(150<X<180)=P(X<180)-P(X\leq150)"

"=P(Z<\dfrac{180-165}{6})-P(Z\leq\dfrac{150-165}{6})"

"=P(Z<2.5)-P(Z\leq-2.5)"

"\approx0.993790-0.006210\approx0.9876"

"98.76\%"


b)


"P(X\geq x)=1-P(Z<\dfrac{x-165}{6})=0.71"

"P(Z<\dfrac{x-165}{6})=0.29"

"\dfrac{x-165}{6}\approx-0.553385"

"x=165-6(0.553385)=162"

71% of the students will receive the score no less than "162"


c)


"P(147<X<159)=P(X<159)-P(X\leq147)"

"=P(Z<\dfrac{159-165}{6})-P(Z\leq\dfrac{147-165}{6})"

"=P(Z<-1)-P(Z\leq-3)"

"\approx0.158655-0.001350\approx0.1573"

"15.73\%"